# Data-Efficient Sleep Staging with Synthetic Time Series Pretraining

### v1

@@ Methods
We based our model architecture on the TinySleepNet architecture, a conceptually simple deep neural network for sleep staging that has previously demonstrated competitive results [21]. This architecture consists of a convolutional feature extractor that extracts features from individual epochs and a classifier that aggregates these feature across multiple epochs to perform sleep staging. Our feature extractor consisted of four convolutional layers with 128 filters each, a kernel size of 50, 8, 8, and 8, respectively, and a stride of 25, 1, 1, and 1, respectively. Following each convolutional layer was a batch normalization layer and a ReLU activation function. The outputs of the first and last convolutional layer were passed through max pooling layers that reduced the temporal dimension of the feature maps by a factor of 8 and 4, respectively. Both max pooling layers were followed by a dropout layer with a dropout rate of 0.5.
The feature extractor was followed by a classifier, which had a different architecture in the pretraining phase than in the fine-tuning phase. During pretraining, the classifier consisted of a dense layer with 80 neurons and ReLU activation function, followed by a dense layer with 20 neurons and sigmoid activation function to predict frequency bins. During fine-tuning, the classifier comprised a bidirectional Long-Short Term Memory (LSTM) [22] layer with a hidden size of 128, a dropout layer with dropout rate 0.5, and a dense layer with 5 neurons and softmax activation function to predict sleep stages. The classifier used for fine-tuning is similar to the original TinySleepNet classifier, except that the unidirectional LSTM layer was replaced with a bidirectional LSTM layer [21].
Pretraining
In the pretraining phase, we generated synthetic time series signals by summing sine waves with random frequencies (see section “Synthetic Data”), and then trained models to identify the frequency bins from which these frequencies were drawn for a given signal (see Fig. 1). Each frequency bin was represented by a single output neuron of the model, with output values greater than 0.5 indicating that the corresponding frequency bin was used to generate the input signal. We trained the models for 20 training epochs using the Adam optimizer [23], a fixed learning rate of $10^{-4}$, a batch size of 64, and a binary cross-entropy loss. This loss function is commonly employed for multi-label classification problems with one-hot encoded labels. It is defined as
@@
where $N$ is the number of samples, $n_{f}$ is the number of frequency bins, $y_{i,j}$ is the true label of sample $i$ for class $j$, and $\hat{y}_{i,j}$ is the predicted label of sample $i$ for class $j$.
@@
where $I$ is the indicator function, $N$ is the number of samples, $n_{f}$ is the number of frequency bins, $y_{i,j}$ is the true label of sample $i$ for class $j$, and $\hat{y}_{i,j}$ is the predicted label of sample $i$ for class $j$.
Fine-Tuning
In the fine-tuning phase, we trained models to predict sleep stages based on sequences of sleep staging data (see Fig. 1). Each sequence (training sample) consisted of 11 epochs, and the feature extractor of our model generated features for each individual epoch. The LSTM-based classifier then aggregated these features and predicted the sleep stage of the middle epoch in the input sequence. To ensure that the model could use the same aggregation process for all epochs of a recording, we padded the first and last five sequences of each recording with zeros to the full sequence length.
@@
We fine-tuned the models using the Adam optimizer [23], a fixed learning rate of $10^{-4}$, weight decay of $10^{-3}$, a batch size of 32, and the categorical cross-entropy loss. To prevent overfitting, we limited each training run to a maximum of 50 training epochs and stopped training early if the loss on the validation data did not improve for 10 training epochs (early stopping). Additionally, we clipped all gradients with a maximum norm greater than 5.0 to prevent exploding gradients.
@@
To track model performance during fine-tuning, we recorded the macro F1 score on the training and validation data after each training epoch. In multi-class settings, the macro F1 score is calculated as the average F1 score [25] across all classes and is defined as
@@
where $k$ is the number of classes, $p_{i}$ is the precision for class $i$, and $r_{i}$ is the recall for class $i$.
Results
Training Configurations
We developed several training configurations to investigate the effectiveness of our pretraining method and the relevance of the learned features for sleep staging. In our experiments, we compared the performance of pretrained models against the performance of non-pretrained models in scenarios with varying amounts of training data. In particular, we studied the performance of our approach in few-samples and few-subject regimes, where the greatest benefit was expected. Furthermore, we analyzed the priors that the model learned during pretraining and the role of frequency information in the learned features. Finally, we investigated whether these features could be further improved by fine-tuning the feature extractor. To answer these questions, we created four training configurations.
Fully Supervised.
The Fully Supervised training configuration is similar to many existing deep learning approaches for sleep staging [3] and served as a baseline to compare our pretrained models against. In this configuration, we skipped the pretraining step and trained (fine-tuned) the model from scratch using sleep staging data.
Fixed Feature Extractor.
We employed the Fixed Feature Extractor configuration to investigate the relevance of the features generated by the pretrained feature extractor for sleep staging. After pretraining the feature extractor, we kept its model weights fixed and only fine-tuned the sleep staging classifier.
Fine-Tuned Feature Extractor.
With this training configuration, we studied (i) how model performance changes when the pretrained feature extractor is allowed to change during fine-tuning and (ii) whether the priors learned during pretraining can prevent overfitting in few-samples or few-subject regimes. As in the previous configuration, we first pretrained the feature extractor, but then fine-tuned the full model without keeping any model weights fixed. Consequently, this configuration is similar to the Fully Supervised configuration with the key distinction that the feature extractor is initialized with pretrained weights.
Untrained Feature Extractor.
The Untrained Feature Extractor configuration was used as a baseline to study whether our pretraining scheme produces priors that are superior to random weights for sleep staging. We randomly initialized the feature extractor using He initialization [26] and then kept its weights fixed while fine-tuning the classifier. This approach mirrors the Fixed Feature Extractor configuration, but with a random feature extractor instead of a pretrained one.
@@ Results
In all four training configurations, we performed training using the cross-validation scheme described in section “Sleep Staging Data”. A single training run in this cross-validation scheme included both pretraining and fine-tuning where applicable. If not specified otherwise, all reported results were obtained on the test folds of the cross-validation splits, while the small part of the training data reserved for validation was used as early stopping set during fine-tuning.
@@
To ensure comparability between different training and data configurations, we seeded the model initialization, the synthetic data generation during pretraining, and the data subsampling when fine-tuning with a limited amount of training data. By using the same seeds for all training configurations, we reduced the influence of random initialization and data sampling on comparisons between training configurations. Furthermore, when training with a reduced amount of data, we maintained a constant number of gradient updates per training epoch by duplicating each training sample $\lfloor\frac{N}{N_{red}}\rfloor$ times, where $N$ is the number of samples in the full training data and $N_{red}$ is the number of samples after reduction.
Data Efficiency
Figure 2: Average macro F1 scores of the different training configurations trained with data from a varying number of subjects. The bars indicate the mean of the macro F1 scores averaged over 15 trainings (3 repetitions of a 5-fold cross-validation) for each training configuration and number of subjects. Error bars show the standard deviation of the macro F1 scores.
To assess the data efficiency of our pretraining method, we trained models from each of the training configurations with the data of a varying number of subjects. The number of subjects in the training data only affected the fine-tuning step, as the pretraining step utilized the same amount of synthetic data regardless of the training data size. A maximum of 56 subjects was available for training, divided into 4 training folds with 14 training subjects (and 2 validation subjects) each. From this pool, we randomly sampled $n_{subj}\in\{1,2,5,10,25,56\}$ subjects to create the training data for a single training run. For each number of subjects $n_{subj}$, we conducted three repetitions of a 5-fold cross-validation scheme, resulting in 15 training runs with 15 macro F1 scores for each training configuration and number of subjects. This approach allowed us to estimate the spread of the macro F1 scores when using different model initializations and data folds for training and testing. Aggregating the scores of multiple training runs was especially important in situations with very little training data, where model performance varied greatly between runs.
Fig. 2 shows the sleep staging performance of the different training configurations (quantified by macro F1 scores) and how this performance depended on the number of subjects included in the training data. The pretrained feature extractors learned features that are more informative for sleep staging than those generated by a random feature extractor. This is especially evident in the low-data regime, where the performance gap between the untrained feature extractor and fixed feature extractor configurations was largest (delta of 0.22 in the average macro F1 score when trained with the data of one subject). Both configurations benefited from an increasing number of subjects in the training data, but the performance gap between the two configurations remained substantial even when trained with the data of all 56 subjects (delta of 0.12 in the average macro F1 score).
When comparing the fixed feature extractor to the fully supervised configuration, our pretraining scheme again appears to be most beneficial in the low-data regime. The fixed feature extractor configuration outperformed the fully supervised configuration by 0.08 in the average macro F1 score when trained with data from only one subject. This performance gap narrowed as more subjects were included in the training data, until both configurations achieved comparable macro F1 scores of 0.67 when trained with data from five subjects. Training with more than five subjects resulted in the fully supervised configuration outperforming the fixed feature extractor configuration.
Fine-tuning the feature extractor after pretraining appeared to combine the advantages of the fixed feature extractor configuration in the low-data regime and the fully supervised configuration in the high-data regime. When trained with the data of only one subject, the fine-tuned feature extractor configuration achieved an average macro F1 score of 0.51, which was close to the performance of the fixed feature extractor configuration (macro F1 score of 0.53) and outperformed the fully supervised configuration (macro F1 score of 0.45). When fine-tuned with the full training data, the fine-tuned feature extractor configuration achieved an average macro F1 score of 0.81, which was on par with the performance of the fully supervised configuration (macro F1 score of 0.80) and outperformed the fixed feature extractor configuration (macro F1 score of 0.76). Overall, the fine-tuned feature extractor configuration achieved similar or better performance than the other training configurations across all numbers of subjects in the training data.
Few-Samples and Few-Subject Regimes
Figure 3: Detailed comparison of the Fully Supervised and Fine-Tuned Feature Extractor configurations in few-samples and few-subject regimes. Panels a and b show the mean and the standard deviation of the macro F1 scores over 15 trainings (3 repetitions of a 5-fold cross-validation) of the fully supervised trained and the fine-tuned feature extractor configurations, respectively. Panel c shows the average macro F1 scores achieved by the fine-tuned feature extractor configuration minus the average macro F1 scores achieved by the fully supervised configuration. These differences were calculated using a bootstrapping approach with 10,000 bootstrap samples. For each bootstrap sample, we first sampled 15 macro F1 scores with replacement from the 15 values available for each matrix entry. Then, we calculated the difference between the averages of the sampled scores. Finally, we display the mean and the standard deviation over these bootstrap samples.
In the previous experiment, we controlled the amount of data available for training subject-wise, which influenced both (i) the total amount of training samples (moving towards the few-samples regime) and (ii) the diversity of the subjects these samples stem from (moving towards the few-subject regime). To disentangle the impact of these two factors on model performance, we carried out a second experiment where we controlled the total number of training samples and the diversity of the subjects in the training data separately. We randomly sampled $n_{subj}\in\{1,2,3,4,5\}$ subjects from the training data and then randomly sampled $n_{samples}\in\{50,130,340,900,all\}$ training samples from the combined data of these subjects (“all” indicates that all available training samples were used). As in the first experiment, we conducted three repetitions of 5-fold cross-validation for each parameter combination. This experiment only focused on the fully supervised and fine-tuned feature extractor training configurations.
Fig. 3 shows the effects that subject diversity and training sample volume had on sleep staging performance. The results demonstrate that both the fully supervised and the fine-tuned feature extractor configurations benefited from an increased subject diversity, even if the total number of training samples was kept constant (see rows in figures 3a and 3b, respectively). Similarly, both configurations benefited from an increased number of training samples when keeping the number of subjects constant (see columns in figures 3a and 3b). Overall, the impact that a reduced subject diversity had on model performance appeared to be comparable to the impact of a reduced number of training samples.
Similar to the observations made in section “Data Efficiency”, the fine-tuned feature extractor configuration generally achieved better macro F1 scores than the fully supervised configuration both in few-samples (see lower part of Fig. 3c) and few-subject regimes (see left part of Fig. 3c). The pretrained models outperformed the fully supervised models by up to 0.11 in macro F1 score. This highlights the potential of our pretraining scheme in situations where only a few subjects and/or annotated epochs are available.
Priors Towards Frequency Information
To get a better understanding of the pretraining process and the priors learned by the model, we recorded several metrics during pretraining. Fig. 4 shows the development of the loss function and the hamming metric during a single pretraining run. Additionally, it shows the accuracy of the fully pretrained model to predict frequencies belonging to each of the 20 frequency bins. The loss function in Fig. 4a converged to a low value, indicating that the model has learned effectively. At the same time, the hamming metric in Fig. 4b reached a high value of 0.9 on the validation data. This value can be interpreted as the model predicting the frequency bins that were used to create the synthetic signals with an accuracy of 90%. Interestingly, Fig. 4c shows that the model was especially proficient in predicting higher frequencies starting from 2.5 Hz (accuracies $>90\%$). Lower frequencies, especially those below 1 Hz, were predicted with lower accuracy (75–85%).
We hypothesize that the differences in prediction accuracy across frequency bins are not due to the pretrained model being unable to predict lower frequencies. Instead, we believe this discrepancy arises from the varying width of the frequency bins (see x-axis of Fig. 4c). Due to their logarithmic scaling, the bins for lower frequencies were narrower than those for higher frequencies. This could potentially make it more challenging for the model to distinguish between the lower frequency bins.
Figure 4: Pretraining metrics of a single pretraining run. Panel a shows the development of the loss function during pretraining for both the training and validation data. The following two panels quantify the accuracy of the model to predict the frequencies of the synthetic signals. The hamming metric in panel b measures the overall accuracy summarized across all frequency bins, while panel c shows the accuracy for each frequency bin separately.
Discussion
In this work, we propose a novel pretraining scheme for EEG time series data that leverages synthetic data generated by a simple random process. We hypothesize that our pretraining scheme could be particularly beneficial in few-samples and few-subject regimes, which we argue could benefit greatly from the priors towards frequency information that a model learns during pretraining.
Our results confirm the effectiveness of our pretraining scheme, particularly in few-samples and few-subject regimes. Pretrained models outperformed fully supervised models when trained with a reduced number of subjects or training samples (see columns and row in Fig. 3c, respectively). This supports observations made in the field of Self-Supervised Learning (SSL) that pretrained models generally have better data efficiency than fully supervised ones [12, 27]. In contrast to SSL methods, however, our pretraining scheme improves data efficiency without requiring empirical data. We hypothesize that this effect is caused by the priors that the model learns during pretraining. These priors could prevent overfitting to a small number of training samples, particularly those from minority classes (e.g., N1), or subject-specific features, which is especially problematic in situations with very little training data. As expected, we observed that all of our training configurations improved with a larger training dataset (see Fig. 2). This aligns with the prevalent view in the literature that deep learning models for sleep staging need substantial amounts of diverse data to perform well [3, 28, 4, 29]. When trained with the full training data, pretrained models performed comparably to fully supervised models (see Fig. 2), achieving macro F1 scores similar to those of other deep learning approaches for sleep staging [3, 30].
We further observed that, while the frequency content of a signal is crucial for sleep staging, deep neural networks are capable of extracting additional information from the data that exceeds the frequency domain. The importance of the frequency content of a signal for sleep staging is demonstrated by the high macro F1 scores achieved by the fixed feature extractor configuration and the substantial performance improvements it achieved over untrained feature extractors (see Fig. 2). We attribute the performance gap between the two training configurations to the priors learned during pretraining the feature extractor. These priors biased the model to extract frequency information from the data, which it achieved with high accuracy after pretraining (see Fig. 4b and c). Our finding is consistent with previous studies that reported frequency-based features to be important for sleep staging [18]. When the feature extractor of our model was allowed to be fine-tuned after pretraining, model performance increased in all data regimes (except for the lowest data regime with only one subject; see Fig. 2). We hypothesize that this increase in the macro F1 score is due to the feature extractor learning to extract information beyond the frequency content of the signal during fine-tuning. This hypothesis is in line with the AASM annotation guidelines [17], which consider several frequency-unrelated features essential for sleep staging. These features include time-domain information, which is important for spindles, as well as amplitudes and specific patterns, such as k-Complexes [4]. Recent studies that applied feature engineering approaches to sleep staging further support our hypothesis by including additional features from the time-domain in their models [31, 32]. In one study, the authors analyzed the most important features for their model and found that time-related features, such as the time elapsed from the beginning of a recording, were among the top 20 most important features [31]. Although it remains unclear what additional information our pretrained models learn during fine-tuning, our method offers a promising avenue for future research into the interpretability of deep neural networks for sleep staging.
There are several opportunities for future work that could build upon our findings. One promising direction is to explore the pretraining task in more detail, for example, by investigating the impact of changing the frequency range that is used to generate the synthetic signals during pretraining. Similar to previous work in the vision domain [15], it could also be promising to investigate which structural properties of synthetic time series are important for sleep staging. This could be achieved by defining new pretraining tasks that are based on different data generation processes and may lead to a better understanding of what constitutes “natural” EEG time series. In addition, we suggest exploring models with greater capacity and less inductive bias, such as transformer models [33, 34], which we expect to benefit even more from our pretraining method. Pretraining such models with synthetic data may alleviate their need for large amounts of training data [35]. Another avenue for future research is to investigate the generalizability of our method to more diverse datasets. In this work, we tested our method only on two datasets with two different pathologies (healthy and diagnosed with sleep apnea). Initiatives such as the sleepdata.org platform [36] and the Temple University data corpus [37] provide a wide range of datasets that could be used for this purpose. Finally, it could be insightful to compare our approach with recent SSL methods [11] and data augmentation strategies that employ synthetic EEG generators [7, 13]. To enable such comparisons and to facilitate future research in this direction, we make our code available online [19].
Our method presents a novel solution to address important issues that affect current deep learning models in the EEG time series domain, without requiring large amounts of patient data. We expect our approach to be advantageous in various applications where EEG data is scarce or derived from a limited number of subjects, such as brain-computer interfaces [2].

### v2

@@ 4.1.2 Model
We based our model on the TinySleepNet architecture, a conceptually simple deep neural network for sleep staging that has previously demonstrated competitive results [41]. This architecture consists of a convolutional feature extractor that extracts features from individual sleep epochs and a classifier that aggregates these feature across multiple epochs to perform sleep staging. The feature extractor consisted of four convolutional layers with 128 filters each, a kernel size of 50, 8, 8, and 8, respectively, and a stride of 25, 1, 1, and 1, respectively. Following each convolutional layer was a batch normalization layer and a ReLU activation function. The outputs of the first and last convolutional layer were passed through max pooling layers that reduced the temporal dimension of the feature maps by a factor of 8 and 4, respectively. After both max pooling layers, we applied dropout with a dropout rate of 0.5.
The feature extractor was followed by a classifier, which consisted of a dense layer with 80 neurons and ReLU activation function as well as a dense layer with 20 neurons and sigmoid activation function (i.e., one output for each frequency bin). We defined a threshold of 0.5 to determine whether a frequency from a frequency bin was present in the signal (output greater than 0.5) or not (output less than or equal to 0.5).
4.1.3 Pretraining for Frequency Prediction
We pretrained models for 20 training epochs using the Adam optimizer [29], a fixed learning rate of $10^{-4}$, a batch size of 64, and a binary cross-entropy loss. This loss function is commonly employed for multi-label classification problems with one-hot encoded labels. It is defined as
@@ 4.1.3 Pretraining for Frequency Prediction
where $N$ is the number of samples, $n_{f}$ is the number of frequency bins, $y_{i,j}$ is the true one-hot encoded label of sample $i$ and frequency bin $j$, and $\hat{y}_{i,j}$ is the predicted one-hot encoded label of sample $i$ and frequency bin $j$.
@@
where $I$ is the indicator function, which assumes the value 1 if the true one-hot encoded label $y_{i,j}$ of sample $i$ and frequency bin $j$ is equal to the predicted one-hot encoded label $\hat{y}_{i,j}$ of sample $i$ and frequency bin $j$.
4.2 Phase 2: Fine-Tuning with Sleep Staging Data
In the fine-tuning phase, we used the pretrained feature extractor of the pretraining phase for sleep stage classification on EEG and EOG data. Sleep stage classification is a multi-class classification problem where a short segment of EEG and EOG data (sleep epoch) is assigned to one of five sleep stages (Wake, N1, N2, N3, REM) based on the underlying brain activity. In our approach, the feature extractor of our model produced features from individual sleep epochs of EEG and EOG data. The extracted features were then concatenated over a sequence of multiple epochs and aggregated by a classifier to predict the sleep stage of the middle epoch in the input sequence (see Figure 1). By incorporating the surrounding sleep epochs into the classifier’s input, we provided additional temporal context, potentially enhancing the accuracy of sleep stage prediction.
4.2.1 Sleep Staging Data
To test our approach, we used four publicly available datasets encompassing a diverse range of subjects, including both healthy individuals and those with various medical conditions such as sleep apnea, REM sleep behavior disorder, and affective disorders: DODO [21], DODH [21], Sleep-EDFx [27, 34], and ISRUC [28]. The DODO and DODH datasets contain 55 recordings from subjects diagnosed with obstructive sleep apnea (OSA) and 25 recordings from healthy subjects, respectively. We combined the two datasets into a single dataset with 80 recordings, which we refer to as DODO/H. Each recording stems from a different subject and was split into 30-s non-overlapping windows (sleep epochs), which were annotated with sleep stages (Wake, N1, N2, N3, REM) by five sleep experts following the AASM guidelines [5]. We aggregated these five expert annotations into a consensus label for each sleep epoch using majority voting, with ties resolved by selecting the sleep stage determined by the most reliable scorer (defined as the scorer with the highest average agreement with all other scorers) [21]. Epochs that the consensus of the sleep experts annotated as artifacts were removed from the datasets. In our experiments, we focused on three EEG and EOG derivations available in both datasets and recommended by the AASM guidelines [5]: C3-M2, F3-M2, EOG1. All channels were sampled at 250 Hz.
The Sleep-EDFx dataset is provided on the PhysioNet platform [17] and contains 197 recordings from two studies [27, 34], of which we used the 153 recordings of the “Sleep Cassette” study [34]. The recordings stem from 78 healthy subjects with two recordings per subject, except for subjects 13, 36, and 52, who have only one recording each due to technical issues. The recordings were split into 30-s sleep epochs and annotated with sleep stages (Wake, N1, N2, N3, N4, REM) according to the 1968 Rechtschaffen and Kales manual [38] by a single expert per recording. To be consistent with the annotations of the other datasets, we combined the N3 and N4 stages into a single N3 stage and removed data annotated as artifacts or movement. Furthermore, we removed excess daytime data from the recordings by cropping them to begin 30 min before the first annotated non-Wake stage and end 30 min after the last annotated non-Wake stage. All available EEG and EOG channels (Fpz-Cz, Pz-Oz, EOG horizontal) were sampled at 100 Hz.
The ISRUC dataset [28] contains 126 recordings and is divided into three subgroups. Subgroup one contains 100 recordings from 100 subjects suffering from various disorders affecting sleep (including OSA, REM disorder, affective disorder, snoring), subgroup two contains 16 recordings from 8 subjects diagnosed with OSA and snoring (2 recordings per subject), and subgroup three contains 10 recordings from 10 healthy subjects. Due to technical issues, we excluded two recordings (recordings 8 and 100) from the first subgroup. All recordings were split into 30-s sleep epochs and annotated with sleep stages by two experts following the AASM guidelines [5]. We only used the annotations of the first expert and focused on the EEG and EOG channels C3-A2, F3-A2, and LOC-A2 sampled at 200 Hz.
To prepare the data for training, we filtered the signals between 0.3 and 35 Hz with an 8th-order zero-phase Butterworth filter and downsampled the filtered signals to 100 Hz using polyphase filtering if necessary. We then normalized the amplitudes of each individual sleep epoch by subtracting the median and dividing the result by the interquartile range of its amplitude distribution. Finally, we clipped all signal amplitudes above 20 or below -20 to minimize outliers.
After preprocessing, the three datasets (DODO/H, Sleep-EDFx, ISRUC) were split subject-wise into five folds for cross-validation. Note that we created a separate cross-validation split for each dataset, resulting in three different 5-fold cross-validation configurations. When splitting the DODO/H and the ISRUC datasets, we ensured that the recordings within each split were balanced in regard to the sub-datasets or subgroups. In addition to splitting the data into training and test folds, we reserved approximately 10% of the data from each fold as validation data for hyperparameter tuning and early stopping (i.e., one recording of DODO and DODH each, two subjects of Sleep-EDFx, and two recordings from the first subgroup of ISRUC for each fold). We validated our models on the validation data of the four training folds. The validation data of the test fold was kept separate and not involved in validation or testing to prevent data leakage.
4.2.2 Model
Since the sleep epochs used for fine-tuning had the same format as the synthetic data generated for pretraining (i.e., 3 channels with 30 s of data sampled at 100 Hz), we kept the architecture of the feature extractor unchanged. However, we replaced the classifier with a different architecture to predict sleep stages instead of frequency bins. The classifier was inspired by the TinySleepNet architecture [41] and consisted of a bidirectional Long-Short Term Memory (LSTM) [25] layer with a hidden size of 128, a dropout layer with a dropout rate of 0.5, and a dense layer with 5 neurons and softmax activation function.
4.2.3 Fine-Tuning for Sleep Stage Classification
@@
We fine-tuned the models with training samples consisting of 11 sleep epochs that provided the classifier with context information from 5 epochs before and after the epoch to be classified. To ensure that the model could use the same aggregation process for all sleep epochs of a recording, we padded the first and last five sequences of each recording with zeros to the full sequence length. Fine-tuning was performed using the Adam optimizer [29], a fixed learning rate of $10^{-4}$, weight decay of $10^{-3}$, a batch size of 32, and the categorical cross-entropy loss. To prevent overfitting, we limited each training run to a maximum of 50 training epochs and stopped training early if the macro F1 score on the validation data did not improve for 10 training epochs (early stopping). Additionally, we clipped all gradients with a maximum norm greater than 5.0 to prevent exploding gradients.
@@
We tracked model performance by recording the macro F1 score on the training and validation data after each training epoch. The macro F1 score is calculated as the average F1 score [42] across all sleep stages and is defined as
@@
where $k$ is the number of sleep stages, $p_{i}$ is the precision of sleep stage $i$, and $r_{i}$ is the recall of sleep stage $i$.
@@
In all four training configurations (Fully Supervised, Fixed Feature Extractor, Fine-Tuned Feature Extractor, and Untrained Feature Extractor), we performed training using the cross-validation scheme described in Section 4.2.1. A single training run in this cross-validation scheme included both pretraining and fine-tuning where applicable. If not specified otherwise, all reported results were obtained on the test folds of the cross-validation splits, while the small part of the training data reserved for validation was used as early stopping set during fine-tuning.
@@
To ensure comparability between different training and data configurations, we seeded the model initialization, the synthetic data generation during pretraining, and the data subsampling when fine-tuning with a limited amount of training data. This strategy reduced the influence of random initialization and data sampling on comparisons between training configurations. In particular, seeding ensured that models were fine-tuned with the same data across all training configurations in the few-sample and few-subject regimes. Furthermore, when training with a reduced amount of data, we maintained a constant number of gradient updates per training epoch by duplicating each training sample $\lfloor\frac{N}{N_{red}}\rfloor$ times, where $N$ is the number of samples in the full training data and $N_{red}$ is the number of samples after reduction.
Experiments were performed on an NVIDIA DGX A100 workstation with eight NVIDIA A100 GPUs and a Dell workstation with a single NVIDIA RTX A6000 GPU.
Author Contributions
Conceptualization, N.G. and S.B.; methodology, N.G.; software, N.G.; validation, N.G.; formal analysis, N.G.; investigation, N.G., S.M. and S.B.; resources, S.B.; data curation, N.G.; writing—original draft preparation, N.G. and S.B.; writing—review and editing, N.G., S.M. and S.B.; visualization, N.G. and S.B.; supervision, S.M. and S.B.; project administration, N.G. and S.B. All authors have read and agreed to the published version of the manuscript.
Funding
This research received no external funding.
Data Availability
The DODO and DODH datasets analyzed in the current study are described in Guillot et al. [21] and are publicly available (see https://github.com/Dreem-Organization/dreem-learning-open (accessed on 30 March 2023) for how to access the datasets). The Sleep-EDFx dataset is described in Kemp et al. [27] and Mourtazaev et al. [34] and is publicly available at https://physionet.org/content/sleep-edfx/1.0.0/ (accessed on 30 March 2023). The ISRUC dataset is described in Khalighi et al. [28] and is publicly available at https://sleeptight.isr.uc.pt/(accessed on 13 April 2023). The source code necessary to reproduce our results is available on GitHub at https://github.com/dslaborg/frequency-pretraining/tree/full-paper-version (accessed on 29 August 2025).